# OminiControl: Minimal and Universal Control for Diffusion Transformer

### v2

@@ 1 Introduction
Diffusion models[9, 28, 25] have revolutionized the field of visual generation, demonstrating remarkable capabilities that significantly outperform traditional approaches like Generative Adversarial Networks (GANs)[6] in terms of image quality and diversity. While these models excel at generating highly realistic imagery, a critical challenge persists: enabling precise and flexible control over the generation process to accommodate diverse and complex user requirements.
Text-based conditioning has been a cornerstone in advancing controllable generation[28, 25, 13, 2, 23, 35], offering an intuitive interface for users to specify their desired outputs. However, text prompts alone often fail to convey precise spatial details and structural attributes that users wish to control. Consequently, recent research has explored complementary conditioning modalities for guiding diffusion models, with image-based control emerging as a particularly effective approach[41, 39, 22, 43, 15]. This multi-modal conditioning strategy enables more detailed and accurate control over the generation process, addressing the limitations inherent in purely text-based interfaces.
Current image conditioning methods can be broadly categorized into spatially aligned and non-spatially aligned approaches. Spatially aligned tasks such as sketch-to-image and inpainting require direct correspondence between conditioning and output images, typically achieved through methods like ControlNet[41] that inject conditioning features in a spatially-preserving manner. In contrast, non-spatially aligned applications including subject driven generation and style transfer, as demonstrated by IP-Adapter[39], often employ pre-trained encoders like CLIP[27] to extract global features for integration via cross-attention mechanisms.
Despite the effectiveness of existing image-conditioned approaches, they present several limitations that hinder their efficiency and flexibility [41, 39, 22]. First, most existing methods are designed specifically for UNet-based architectures [29, 41, 22, 39, 31, 42, 43, 40, 24, 16], as seen in Stable Diffusion models [28, 25]. While these approaches work well with UNet’s encoder-decoder structure, they may not translate effectively to the more advanced Diffusion Transformer (DiT) models[23] that have demonstrated superior image generation quality[13, 2]. Additionally, current approaches typically specialize in either spatially aligned [22, 43, 41] or non-spatially aligned tasks [39, 42, 17, 15, 12], lacking a unified architecture to handle both control types effectively. This specialization often requires practitioners to employ different methods for different control scenarios, increasing system complexity and implementation overhead. Furthermore, these methods rely heavily on additional network structures [39, 42, 22, 43, 41, 17], which introduce substantial parameter overhead.
To address these limitations, we propose a parameter-efficient approach for incorporating image-based control into DiT architectures[23]. Our method reuse the model’s existing VAE encoder[28] to process conditioning images. Following the same token processing pipeline as noisy image tokens, we augment the encoded features with learnable position embeddings[34], and seamlessly integrate them alongside latent noise in the denoising network. This design enables direct multi-modal attention interactions[23, 30] between condition and generation tokens throughout the DiT’s transformer blocks, facilitating efficient information exchange and control signal propagation.
@@
We implemented our method on the high-performing DiT-structured diffusion model, FLUX.1-dev[13], a large-scale model containing 12 billion parameters. Extensive experiments on edge-guided generation, depth-aware synthesis, region-specific editing, and identity-preserving generation indicate that our DiT-based approach yields better results compared to both UNet-based implementations[41, 39, 7] and their community adaptations on the FLUX.1 model [14, 37].
For identity-preserving generation, we developed a novel data synthesis pipeline that generates high-quality, identity-consistent image pairs. Using this pipeline, we created a comprehensive dataset comprising over 200,000 diverse images. To facilitate future research in this direction, we will release both our dataset and the complete pipeline implementation as open-source resources [^2] More details are available in the supplementary material..
In summary, we highlight our contributions as follows:
  1. 1.
    We present a parameter-efficient method for enabling image-conditioned control in Diffusion Transformer (DiT) models, achieving both spatially aligned and non-spatially aligned control within a unified framework.
  2. 2.
    We demonstrate the effectiveness of our approach through extensive experiments across diverse control tasks, including edge-guided generation, depth-aware synthesis, region-specific editing, and identity-preserving generation, consistently outperforming existing methods on both UNet implementations and their DiT adaptations.
  3. 3.
    We develop and release Subjects200K, a high-quality dataset of over 200,000 subject-consistent images, along with an efficient data synthesis pipeline, providing valuable resources to the research community for further exploration of subject-consistent generation tasks.
@@ 2 Related works
2.1 Diffusion-based models
Diffusion-based methods have emerged as a powerful framework for image generation[9, 28], demonstrating success across diverse tasks including text-to-image synthesis [28, 2, 33], image-to-image translation [32], and image editing [20, 1]. Recent advances have led to significant improvements in both quality and efficiency, notably through the introduction of latent diffusion models [28]. To further enhance generative capabilities, large-scale transformer architectures have been integrated into these frameworks, leading to advanced models like DiT[23, 2, 13, 3]. Building on these architectural innovations, FLUX[13] incorporates transformer-based design with flow matching objectives [18], achieving state-of-the-art generation performance.
@@ 2.2 Controllable generation with diffusion models
Controllable generation has been extensively studied in the context of diffusion models. Text-to-image models[28, 25] have established a foundation for conditional generation, while various approaches have been developed to incorporate additional control signals such as image. Notable methods include ControlNet [41], enabling spatially aligned control in diffusion models, and T2I-Adapter [22], which improves efficiency with lightweight adapters. UniControl [26] uses Mixture-of-Experts (MoE) to unify different spatial conditions, further reducing model size. However, these methods rely on spatially adding condition information to the denoising network’s hidden states, inherently limiting their effectiveness for non-spatial tasks like subject-driven generation. IP-Adapter [39] addresses this by introducing cross-attention through an additional encoder, and SSR-Encoder [42] further enhances identity preservation in image-conditioned tasks. Despite these advances [15, 19, 5], a unified solution for both spatially aligned and non-aligned tasks remains elusive.
@@ 3.1 Preliminary
The Diffusion Transformer (DiT) model [23], employed in architectures like FLUX.1 [13], Stable Diffusion 3 [28], and PixArt [2], uses a denoising network of transformer blocks to refine noisy image tokens iteratively.
Each transformer block processes two types of tokens: noisy image tokens $\mathbf{X}\in\mathbb{R}^{N\times d}$ and text condition tokens $\mathbf{C}_{\text{T}}\in\mathbb{R}^{M\times d}$, where $d$ is the embedding dimension, $N$ and $M$ are the number of image and text tokens respectively (Figure 2). These tokens are embedded into hidden states $X$ and $C_{\text{T}}$, which maintain consistent shapes throughout the transformer blocks.
In each DiT block, after normalizing $X$ and $C_{\text{T}}$, they are processed by the core MM-Attention module [30], which employs Rotary Position Embedding (RoPE) [34] to incorporate positional dependencies across tokens. For a token at position $(i,j)$ in the 2D grid, RoPE applies rotation matrices to the query and key projections:
$Q_{X}(i,j)$ $=W_{Q}\left(X_{i,j}\cdot R(i,j)\right),$ (1)
$K_{X}(i,j)$ $=W_{K}\left(X_{i,j}\cdot R(i,j)\right),$ (2)
where $R(i,j)$ is the rotation matrix at position $(i,j)$. Similarly, the text condition tokens $C_{\text{T}}$ have their query and key projections defined in the same way, with all text token positions set to $(0,0)$ in FLUX.1.
After applying RoPE, the queries, keys, and values from both token types are concatenated to form unified matrices $Q_{Z}$, $K_{Z}$, and $V_{Z}$, representing the combined token set $Z=[X;C_{\text{T}}]$. The MM-Attention operation is then computed as:
$\text{MMAttention}(Z)=\text{softmax}\left(\frac{Q_{Z}K_{Z}^{\top}}{\sqrt{d}}\right)V_{Z},$ (3)
enabling interactions between image and text condition tokens through the attention mechanism.
@@
3.2 Image condition integration
Our approach first encodes the condition image through the model’s VAE, projecting it into the same latent space as the noisy image tokens to form $\mathbf{C}_{\text{I}}\in\mathbb{R}^{N\times d}$.
Previous methods like ControlNet [41] and T2I-Adapter [22] incorporate the condition image by spatially aligning and adding its hidden states directly to those of the noisy image tokens:
$H_{X}=H_{X}+H_{C_{\text{I}}},$ (4)
where $H_{X}$ represents the combined hidden states for further processing, with $H_{C_{\text{I}}}$ being the hidden states from the condition image. While this approach proves effective for spatially aligned tasks, it faces two key limitations: (1) it lacks flexibility when handling non-aligned scenarios, and (2) even in spatially aligned cases, the direct addition of hidden states constrains token interactions, potentially limiting the model’s performance.
@@ 3.2 Image condition integration
In contrast, to enable non-aligned control tasks and provide greater conditioning flexibility, our method processes condition image tokens uniformly with text and noisy image tokens, integrating them into a unified sequence:
$Z=[X;C_{\text{T}};C_{\text{I}}],$ (5)
where $Z$ represents the concatenated sequence of noisy image tokens $X$, text tokens $C_{\text{T}}$, and condition image tokens $C_{\text{I}}$. This unified approach enables direct participation in multi-modal attention [30] without specialized processing pathways (illustrated in Figure 2).
@@
The comparative result shows that our approach achieves higher generation quality and better alignment with the conditions compared to the direct adding method, as illustrated in Figure 3. Moreover, the training curves in Figure 4(a) demonstrate that multi-modal attention method consistently achieves lower loss values than the direct adding approach. Besides, the effectiveness of this unified sequence approach is demonstrated across both spatially aligned and non-spatially aligned tasks (Figure 5), highlighting its versatility in handling diverse conditional generation scenarios.
3.3 Adaptive position embedding
Our unified sequence design allows for flexible integration of condition image tokens, but this requires incorporating positional information to ensure effective interaction with noisy image tokens. The relative positioning of these tokens is critical, as it directly affects the model’s learning efficiency and generalization capability.
In FLUX.1’s Transformers, each token is assigned a corresponding position index to encode spatial information. For a 512$\times$512 target image, the VAE [11] encoder first projects it into the latent space, then the latent representation is divided into a 32$\times$32 grid of tokens, where each token is assigned a unique two-dimensional position index $(i,j)$ with $i,j\in[0,31]$. This indexing scheme preserves the spatial structure of the original image in the latent space, while text tokens maintain a fixed position index of $(0,0)$.
@@ 3.3 Adaptive position embedding
For spatially aligned tasks, our initial approach was to assign condition tokens the same position embeddings as their corresponding tokens in the noisy image. However, for non-spatially aligned tasks such as subject-driven generation, our experiments revealed that shifting the position indices of condition tokens leads to faster convergence (Figure 4(b)). Specifically, we shift the condition image tokens to indices $(i,j)$ where $i\in[0,31]$ and $j\in[32,64]$, ensuring no spatial overlap with the original image tokens $X$.
3.4 Condition strength factor
The unified attention mechanism we adopt not only enables flexible token interaction but also allows us to precisely control the influence of condition images. Specifically, we designed a method that allows for manual adjustment of the condition image’s effect during inference. For a given strength factor $\gamma$, setting $\gamma=0$ removes the condition image’s influence, resulting in an output based purely on the original input. At $\gamma=1$, the output fully reflects the condition image, and as $\gamma$ increases beyond 1, the condition’s effect becomes even more pronounced.
@@ 3.4 Condition strength factor
To achieve this controllability, we introduce a bias term into the original MM-Attention operation. Specifically, we modify Equation 3 to:
$\text{BiasedAttention}(Z)=\text{softmax}\left(\frac{Q_{Z}K_{Z}^{\top}}{\sqrt{d}}+\text{bias}(\gamma)\right)V_{Z},$ (6)

### v4

@@ 1 Introduction
The ability to generate high-quality images with precise user control remains a central challenge in computer vision. While diffusion models [10, 28] have significantly advanced the state of the art in image generation, surpassing traditional GAN-based approaches [7] in both quality and diversity, the problem of fine-grained control persists. Text-conditioned diffusion models [28, 25, 5, 14] serve as the primary paradigm for controllable generation, enabling users to guide the synthesis process through natural language descriptions. However, these models fundamentally lack the capacity to specify exact spatial details and visual attributes that users often need to control.
To address this limitation, recent works have turned to image-based control methods [42, 41, 22, 44, 16, 34, 38, 17, 26, 37] as a promising solution to achieve fine-grained control. These approaches enable users to specify their intentions through reference images or visual hints, offering more precise guidance than text alone [28].
However, current image control methods face several key challenges: First, existing methods require substantial architectural modifications with dedicated control modules, leading to significant parameter overhead [42, 41]. Second, these approaches show clear task bias – they typically work well for either spatially aligned controls (e.g., edge-guided or depth-guided generation [42, 22, 44, 26]) or spatially unaligned ones (e.g., style transfer or subject-driven generation [41, 17, 16, 34]), but rarely both. Third, current approaches are built primarily for UNet architectures [29]. These approaches may yield suboptimal results when applied to newer Diffusion Transformer (DiT) models [24] due to fundamental architectural differences, despite the latter’s superior generation capabilities [14, 2, 5].
These challenges, particularly within the emerging paradigm of Diffusion Transformer models, motivate us to rethink the fundamental approach to image control. We propose OminiControl, an omni-capable yet minimal framework that achieves effective control by leveraging existing components within the DiT architecture.
To achieve the goal of minimal architectural and parameters overhead, OminiControl first utilizes the built-in VAE encoder to project conditioning images into the same latent space as the target images. The conditional tokens are concatenated with the noisy latent and processed by the DiT using multi-modal attention [23, 30, 24]. This unified sequence allows the denoising network to handle both signals directly, requiring only minimal LoRA [4] fine-tuning of the existing transformer blocks. Unlike existing methods such as IP-Adapter [41] that relies on separate CLIP [27] feature extractors and ControlNet [15] that requires additional control modules, OminiControl significantly reduces architectural complexity and parameter overhead.
Building on this, OminiControl unifies controlled generation tasks using the DiT architecture. By modifying explicit position embeddings [35] in the attention mechanism, it handles both spatially aligned and non-aligned tasks. By explicitly controlling condition token positions in the attention mechanism, OminiControl not only handles non-aligned control tasks through semantic attention, but also achieves spatially-aligned control through flexible token interactions, rather than relying on rigid spatial feature addition as in ControlNet [42].
Alongside the design of OminiControl, to fully realize its potential in diverse control scenarios, we address a critical challenge in subject-driven generation – the scarcity of high-quality training data for identity preservation. To this end, we develop an automated data synthesis pipeline specifically designed for subject-driven generation tasks. This pipeline builds upon a key observation that DiT models can generate identity-consistent image pairs with remarkable consistency by simply providing appropriate prompts [14]. Using this pipeline, we create a comprehensive dataset Subjects200K including over 200,000 diverse images with rich identity variations. To facilitate future research, we have released both our dataset and the complete pipeline implementation as open-source resources [^2] Code and dataset are available at: https://github.com/Yuanshi9815/Subjects200K.
Extensive experiments demonstrate that our approach achieves better performance across diverse control tasks. Building upon the state-of-the-art DiT model Flux [14], we evaluate OminiControl on a comprehensive suite of tasks including edge-guided generation, depth-guided generation, and subject-driven generation. Our method outperforms existing approaches such as ControlNet [42], T2I-Adapter [22], and IP-Adapter [41] in their original UNet implementations, as well as their community adaptations on the DiT architecture [15, 39]. Notably, OminiControl achieves these superior results while maintaining remarkable efficiency – requiring only $0.1\%$ additional parameters compared to the base model.
In summary, we highlight our contributions as follows:
  • We propose OminiControl, a minimal and universal control framework for DiT models that: (1) minimizes architectural changes by reusing existing components , introducing only 0.1% additional parameters; and (2) achieves unified control by integrating condition tokens into the original multi-modal attention and leveraging position embeddings to handle both spatially-aligned and unaligned tasks.
  • We develop an automated data synthesis pipeline that leverages DiT’s inherent capability in generating identity-consistent image pairs. Using this, we create and release Subjects200K, a large-scale dataset containing over 200,000 diverse images for subject-driven generation.
  • We conduct extensive experiments on various control tasks including edge-guided, depth-guided, and subject-driven generation. Building upon Flux, OminiControl outperforms existing approaches in both their original UNet implementations and DiT adaptations.
@@ 2 Related works
2.1 Diffusion models
Diffusion models have emerged as a powerful framework for image generation[10, 28], demonstrating success across diverse tasks including text-to-image synthesis [28, 2, 33], image-to-image translation [32], and image editing [20, 1]. Recent advances have led to significant improvements in both quality and efficiency, notably through the introduction of latent diffusion models [28]. To further enhance generative capabilities, large-scale transformer architectures have been integrated into these frameworks, leading to advanced models like DiT[24, 2, 14, 3]. Building on these architectural innovations, FLUX[14] incorporates transformer-based design with flow matching objectives [18], achieving state-of-the-art generation performance.
@@ 2.2 Controllable generation with diffusion models
Controllable generation has been extensively studied in the context of diffusion models. Text-to-image models[28, 25] have established a foundation for conditional generation, while various approaches have been developed to incorporate additional control signals such as image. Notable methods include ControlNet [42], enabling spatially aligned control in diffusion models, and T2I-Adapter [22], which improves efficiency with lightweight adapters. UniControl [26] uses Mixture-of-Experts (MoE) to unify different spatial conditions, further reducing model size. However, these methods rely on spatially adding condition features to the denoising network’s hidden states, inherently limiting their effectiveness for spatially non-aligned tasks like subject-driven generation. IP-Adapter [41] addresses this by introducing cross-attention through an additional encoder, and SSR-Encoder [43] further enhances identity preservation in image-conditioned tasks. Despite these advances [16, 19, 6], a unified solution for both spatially aligned and non-aligned tasks remains elusive.
@@ 3.1 Preliminary
The Diffusion Transformer (DiT) model [24], employed in architectures like FLUX.1 [14], Stable Diffusion 3 [28], and PixArt [2], uses transformer as denoising network to refine noisy image tokens iteratively.
A DiT model processes two types of tokens: noisy image tokens ${X}\in\mathbb{R}^{N\times d}$ and text condition tokens ${C}_{\text{T}}\in\mathbb{R}^{M\times d}$, where $d$ is the embedding dimension, $N$ and $M$ are the number of image and text tokens respectively (Figure 2). Throughout the network, these tokens maintain consistent shapes as they pass through multiple transformer blocks.
In FLUX.1, each DiT block consists of layer normalization followed by Multi-Modal Attention (MMA) [23], which incorporates Rotary Position Embedding (RoPE) [35] to encode spatial information. For image tokens ${X}$, RoPE applies rotation matrices based on the token’s position $(i,j)$ in the 2D grid:
${X}_{i,j}\rightarrow{X}_{i,j}\cdot R(i,j),$ (1)
where $R(i,j)$ is the rotation matrix at position $(i,j)$. Text tokens ${C}_{\text{T}}$ undergo the same transformation with their positions set to $(0,0)$.
The multi-modal attention mechanism then projects the position-encoded tokens into query $Q$, key $K$, and value $V$ representations. It enables the computation of attention between all tokens:
$\text{MMA}([{X};{C}_{\text{T}}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$ (2)
where $[{X};{C}_{\text{T}}]$ denotes the concatenation of image and text tokens. This formulation enables bidirectional attention.
@@
3.2 OminiControl
@@
Building upon the DiT architecture, with FLUX.1 [14] as our implementation base,we aim to develop OminiControl, an omni-capable yet minimal control framework that guides generation through visual conditions.
Leveraging existing architecture
To achieve both omni-capability and minimal architectural changes, OminiControl first reuses the VAE [12, 28] encoder from the base DiT model to project conditioning images into the same latent space as noisy image tokens. This approach contrasts with previous methods that rely on separate feature extractors like CLIP [27, 41] or additional control modules [43], significantly reducing architectural complexity.
The encoded condition tokens ${C}_{\text{I}}$ share the same dimensionality and latent space as the noisy image tokens ${X}$, enabling them to be directly processed by the transformer blocks. Since both condition and image tokens reside in the same latent space, OminiControl leverages the existing DiT blocks to process them jointly, requiring only minimal fine-tuning through LoRA [4] to adapt these blocks for handling the additional condition tokens, instead of duplicating the entire network as done in ControlNet [42].
Unified sequence processing
Building upon these shared latent space, we next need to integrate condition tokens into the model for flexible control.
Previous methods like ControlNet [42] and T2I-Adapter [22] incorporate condition images through direct feature addition:
${X}\leftarrow{X}+{C}_{\text{I}},$ (3)
where the condition features ${C}_{\text{I}}$ are spatially aligned and added to the noisy image tokens ${X}$. While effective for spatially aligned tasks, this approach faces two limitations: (1) it lacks flexibility for non-aligned scenarios where spatial correspondence doesn’t exist, and (2) the rigid addition operation constrains potential interactions between condition and image tokens.
Instead, OminiControl directly concatenates condition tokens with noisy image tokens $[{X};{C}_{\text{T}};{C}_{\text{I}}]$ for multi-modal attention processing. This formulation enables flexible token interactions through DiT’s multi-modal attention mechanism, allowing direct relationships to emerge between any pair of tokens without imposing rigid spatial constraints.
As shown in Figure 5, this approach effectively handles both spatially aligned and non-aligned tasks, with attention maps revealing clear cross-token relationships and interaction patterns. Empirically, this unified sequence approach consistently achieves lower training loss compared to direct feature addition (Figure 4(a)), demonstrating its superior conditioning capability across diverse generation scenarios.
@@
Position-aware token interaction
While the unified sequence approach allows flexible token interactions, encoding position information for the newly appended conditional tokens is not straightforward.
In FLUX.1, the RoPE [35] mechanism requires each token to have an assigned position index $(i,j)$, which might directly influence how tokens attend to each other during multi-modal attention.
Specifically, for a 512$\times$512 input image, the VAE encoder produces a 32$\times$32 grid of latent tokens, with each token assigned a position index $(i,j)$ where $i,j\in[0,31]$. Text tokens ${C}_{\text{T}}$ maintain fixed positions at $(0,0)$ following FLUX.1’s convention. Therefore, it is essential to assign proper position indexing the conditioned tokens.
In our method, we assign position indices differently for spatially aligned and non-aligned tasks. For condition image tokens ${C}_{\text{I}}$, we propose a dynamic positioning strategy based on the control task:
$(i,j)_{{C}_{\text{I}}}=\begin{cases}(i,j)_{{X}}&\text{for aligned tasks}\\ (i,j)+\Delta&\text{for non-aligned tasks}\end{cases}$ (4)
where $\Delta$ is a fixed offset vector.
@@
For spatially aligned tasks like edge-guided generation, condition tokens share positions with their corresponding noisy image tokens, facilitating precise spatial correspondence. For non-aligned tasks like subject-driven generation, we shift condition tokens’ positions by $\Delta$, ensuring no spatial overlap with ${X}$. As demonstrated in Figure 4(b), this shifting leads to faster convergence and better performance for non-aligned tasks.
Controllable conditioning strength
While OminiControl’s unified sequence processing and multi-modal attention enable effective token interactions during training, practical applications often require adjustable conditioning strength at test time.
@@
We achieve this by introducing a bias term into the multi-modal attention computation. Specifically, for a given strength factor $\gamma$, we modify the attention operation in Equation 2 to:
$\text{MMA}([{C}_{\text{T}};{X};{C}_{\text{I}}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+B(\gamma)\right)V,$ (5)
where $B(\gamma)$ is a bias matrix modulating the attention between concatenated tokens $[{C}_{\text{T}};{X};{C}_{\text{I}}]$. Given ${C}_{\text{T}}\in\mathbb{R}^{M\times d}$ and ${X},{C}_{\text{I}}\in\mathbb{R}^{N\times d}$, the bias matrix has the structure:
$B(\gamma)=\begin{bmatrix}\mathbf{0}_{M\times M}&\mathbf{0}_{M\times N}&\mathbf{0}_{M\times N}\\ \mathbf{0}_{N\times M}&\mathbf{0}_{N\times N}&\log(\gamma)\mathbf{1}_{N\times N}\\ \mathbf{0}_{N\times M}&\log(\gamma)\mathbf{1}_{N\times N}&\mathbf{0}_{N\times N}\end{bmatrix}.$ (6)
This formulation preserves the original attention patterns within each token type while scaling attention weights between ${X}$ and ${C}_{\text{I}}$ by $\log(\gamma)$. At test time, setting $\gamma=0$ removes the condition’s influence, while $\gamma>1$ enhances it. This allows users to flexibly adjust the balance between text and image conditions without retraining the model.
3.3 Subjects200K datasets
Figure 6: Examples from our Subjects200Kdataset. Each pair of images shows the same object in varying positions, angles, and lighting conditions. The dataset includes diverse objects such as clothing, furniture, and animals, totaling over 200,000 images.
A critical challenge in developing universal control frameworks lies in obtaining high-quality training data for subject-driven generation. While OminiControl’s architecture enables flexible control, effective training requires data that maintains subject consistency while incorporating natural variations in pose, lighting, and context.
Existing solutions often rely on identical image pairs [41] or limited-scale datasets [31, 13], which present several limitations. Using identical pairs can lead to overfitting in our framework, causing the model to simply reproduce the input. Meanwhile, existing datasets with natural variations often lack either the scale or diversity needed for robust training.